OpenStudy (anonymous):

x^2 + 6x + 11 can be written as (x + h)^2 + k. Find k.

OpenStudy (anonymous):

x^2+6x+11 = x^2+2*3x+9+2=x^2+2*3x+3^2+2=(x+3)^2+2

OpenStudy (anonymous):

so it's clear??

OpenStudy (anonymous):

huh? k is (x+3)^2+2??